# Anant-Net: Breaking the Curse of Dimensionality with Scalable and Interpretable Neural Surrogate for High-Dimensional PDEs

### v2

@@ Abstract
High-dimensional partial differential equations (PDEs) arise in diverse scientific and engineering applications but remain computationally intractable due to the curse of dimensionality. Traditional numerical methods struggle with the exponential growth in computational complexity, particularly on hypercubic domains, where the number of required collocation points increases rapidly with dimensionality. Here, we introduce Anant-Net, an efficient neural surrogate that overcomes this challenge, enabling the solution of PDEs in high dimensions. Unlike hyperspheres, where the internal volume diminishes as dimensionality increases, hypercubes retain or expand their volume (for unit or larger length), making high-dimensional computations significantly more demanding. Anant-Net efficiently incorporates high-dimensional boundary conditions and minimizes the PDE residual at high-dimensional collocation points. To enhance interpretability, we integrate Kolmogorov–Arnold networks into the Anant-Net architecture. We benchmark Anant-Net’s performance on several linear and nonlinear high-dimensional equations, including the Poisson, Sine-Gordon, and Allen-Cahn equations, demonstrating high accuracy and robustness across randomly sampled test points from high-dimensional space. Importantly, Anant-Net achieves these results with remarkable efficiency, solving 300-dimensional problems on a single GPU within a few hours. We also compare Anant-Net’s results for accuracy and runtime with other state-of-the-art methods. Our findings establish Anant-Net as an accurate, interpretable, and scalable framework for efficiently solving high-dimensional PDEs.
@@ 1 Introduction
Physics-informed deep learning (PIDL) represents a rapidly advancing framework that integrates known governing physical laws, typically formulated as PDEs, into the training process of deep neural networks. In contrast to conventional data-driven models that rely solely on observational data, PIDL incorporates physical constraints to guide learning, thereby enhancing generalization, reducing data dependence, and improving interpretability. This synthesis of physics and deep learning has demonstrated broad applicability in solving forward and inverse problems across scientific and engineering domains, particularly in scenarios involving limited, noisy, or deceptive data. Key methodologies under the PIDL umbrella include physics-informed neural networks (PINNs) [1, 2, 3, 4], which embed PDE constraints via automatic differentiation; sparse identification of nonlinear dynamics (SINDy) [5, 6], which infers governing equations by promoting sparsity in learned representations; and neural operators [7, 8, 9, 10, 11], which approximate solution operators across function spaces to model families of PDEs. These approaches are particularly well-suited for high-dimensional problems, where traditional numerical solvers suffer from the curse of dimensionality.
High-dimensional PDEs are integral to various scientific and engineering domains, including quantum mechanics, financial mathematics, and optimal control. Their solutions provide crucial insights into complex, multi-scale phenomena that cannot be accurately captured using lower-dimensional approximations. However, solving these equations efficiently remains a significant challenge due to the curse of dimensionality, the exponential growth in computational complexity and data requirements as the number of dimensions increases. This issue is particularly pronounced in numerical approximations, making high-dimensional PDEs computationally intractable with conventional methods. In this work, we introduce Anant-Net (with Anant meaning infinite), a novel neural architecture designed to mitigate the curse of dimensionality and enable the efficient solution of high-dimensional PDEs critical to real-world applications. Our approach provides an efficient framework for handling large datasets, whose size typically scales exponentially with dimensionality. While purely data-driven models might seem a viable alternative, they often struggle with excessive data requirements, making them impractical for high-dimensional problems. To address this, we leverage physics-informed neural surrogates, incorporating the underlying PDE as an additional constraint. This integration significantly reduces the dependency on large datasets, offering a more scalable and computationally efficient framework for solving high-dimensional PDEs.
@@ 1.1 Related Work
In the context of high-dimensional PDEs, two primary factors determine the efficacy of any method based on scientific machine learning: (1) the sampling strategy for data or collocation points, and (2) the computation of gradients for residual losses via automatic differentiation, which are used as soft constraints in training physics-informed models. It is evident that a significant number of existing methods represent a trade-off between these factors, which ultimately influences their scalability and suitability for solving real-world problems of practical relevance. While various methods exist for solving high-dimensional PDEs, this work focuses on neural surrogates that not only leverage available data but also incorporate physical laws, ensuring the consistency of the solutions with underlying governing principles. In recent years, the challenge of solving high-dimensional PDEs has been addressed by several researchers through the use of neural surrogate methods. Han et al. [12] introduced a deep neural network framework that reformulates high-dimensional PDEs as backward stochastic differential equations. However, their approach relied on approximating the gradient of the solution using a neural network, which introduced scalability issues, particularly in scenarios involving long-time integration. This limitation was subsequently addressed by Raissi et al. [13], who proposed an alternative method that directly models the solution of the PDE with a neural network, rather than its gradient. The efficacy of this improvement was demonstrated through its application to the Black-Scholes-Barenblatt and Hamilton-Jacobi-Bellman equations in 100 dimensions. Hu et al. [14] introduced a stochastic dimensions gradient descent (SDGD) framework based on physics-informed neural networks to solve high-dimensional PDEs on hypercubes. This method stochastically samples the dimensions of the PDE and computes the residuals only in the selected sample dimensions, rather than performing the computation across all dimensions. Their approach significantly reduced the computational burden on hardware, leading to notable speed-ups and improved accuracy compared to other contemporary methods for high-dimensional PDEs. However, their method bypassed the need for data or boundary point sampling by directly enforcing boundary conditions within the model, particularly for hypersphere, where the internal volume approaches zero as the number of dimensions increases (see Figure 1). In the similar line of work, Hu et al.[15] proposed Monte Carlo fractional PINN (MC-fPINN) and its extension, MC-tfPINN for tempered fractional PINN (an extension to fPINN [16]) to solve extremely high-dimensional problems. Further, an alternate approach was proposed by the same authors [17] extended fractional PINN to solve high-dimensional Fokker-Planck-Lévy (FPL) equations in high-dimensions. Adaptive training methods have also been proposed as supplementary strategies to enhance the accuracy of physics-informed neural network models for high-dimensional PDEs [18].
@@
Several studies have leveraged the weak formulation of PDEs to train deep learning models, particularly for high-dimensional problems. A pioneering approach in this direction is the Deep Ritz method [25], which utilizes the variational formulation of PDEs by minimizing an energy functional. This method, designed to address high-dimensional PDEs, employs a ResNet-based architecture and is trained using stochastic gradient descent. It systematically samples data from both the interior and the boundary of the domain, distinguishing itself from approaches such as that of Hu et al. [14], which circumvent boundary data sampling. Building on this framework, the Weak Adversarial Network (WAN) [26] was introduced to solve high-dimensional PDEs by reformulating the problem of finding a weak solution as an operator norm minimization task. In contrast to conventional numerical methods such as finite difference schemes, WAN demonstrated computational efficiency, stability, and the ability to mitigate the curse of dimensionality. The methodology employs generative adversarial network (GAN) framework, where the weak solution and test functions are parameterized by deep neural networks, termed the primal and adversarial networks, respectively. More recently, the WAN framework has been extended through the introduction of XNODE-WAN [27], which replaces the primal network with a Neural Ordinary Differential Equation (Neural ODE) model. This enhancement has resulted in improved computational speed and accuracy, particularly for time-dependent high-dimensional PDEs. These approaches systematically sample data across both the interior and the boundary of the computational domain to enhance the robustness of their solutions.
@@ 1.2 Our Contributions
  • We also introduce an interpretable variant of Anant-Net based on the Kolmogorov–Arnold Network, called as Anant-KAN, enabling further insights into the learned solution structures in high-dimensional settings.
  • We benchmark Anant-Net against several state-of-the-art methods, reporting comparative results in terms of runtime as well as the mean and variance of relative testing error (%), and observe that Anant-Net consistently outperforms the alternatives.
The paper is organized as follows: Section 2 presents a detailed description of the Anant-Net architecture, including the methodologies employed for structuring and integrating training data and collocation points. Section 3 provides numerical results for a representative set of high-dimensional boundary value problems, along with a comparative analysis against existing state-of-the-art architectures. Finally, Section 4 highlights the key findings of this work with a detailed discussion.
@@ 2 Methodology: Anant-Net Architecture
Figure 2: Schematic representation of general Anant-Net architecture with batch-size $B$, highlighting its core components and computational flow for solving high-dimensional PDEs.
@@
Figure 2 shows the general architecture of Anant-Net. The architecture comprises a number of body networks equal to the batch size. Let $u\in\mathbb{R}^{d}$ denote the solution to a $d$-dimensional PDE, representing the unknown quantity to be approximated by Anant-Net. For brevity, we consider a batch size of three, the Anant-Net prediction of the solution field, denoted by $\hat{u}_{\theta}$ , is expressed as:
$\hat{u}_{\theta}(X_{1},X_{2},X_{3})=f_{1}^{(\theta_{1})}(X_{1})\otimes f_{2}^{(\theta_{2})}(X_{2})\otimes f_{3}^{(\theta_{3})}(X_{3}),$ (1)
where the inputs are partitioned as $X_{1}=\{x_{1},\dots,x_{d/3}\}$, $X_{2}=\{x_{d/3+1},\dots,x_{2d/3}\}$ , and $X_{3}=\{x_{2d/3+1},\dots,x_{d}\}$, such that $X_{1}\cap X_{2}\cap X_{3}=\emptyset$. The indices $p\in\mathrm{D}_{1}=\{1,2,\ldots,d/3\},q\in\mathrm{D}_{2}=\{d/3+1,2,\ldots,2d/3\},r\in\mathrm{D}_{3}=\{2d/3+1,2,\ldots,d\}$ , corresponding to the number of batch elements, are stochastically sampled from the set of active dimensions, i.e., $\{x_{p},x_{q},x_{r}\}\subset X=X_{1}\cup X_{2}\cup X_{3}$, which sweeps over all dimensional indices. The network parameters are given by $\theta=\bigcup\theta_{i},\ \forall i$ (all body networks), where $f_{j}^{(\theta_{i})}$ denotes the output vector of the $i^{\text{th}}$ body network. Although there is no strict ordering required for partitioning the dimension set $\mathrm{D}=\mathrm{D}_{1}\cup\mathrm{D}_{2}\cup\mathrm{D}_{3}$, we adopt an ascending-order partitioning strategy for all experiments presented in this paper for the sake of consistency. The number of partitioned subsets is arbitrary and can be increased by proportionally increasing the number of body networks in the Anant-Net architecture. However, for brevity, we fix the total number of partitioned sets; and correspondingly, the number of body networks, to three in all experiments discussed herein. Algorithm 1 gives the details of training data and collocation points for the Anant-Net architecture. It is also important to note that modifying the number of body networks by changing the number of partitioned dimension subsets must be reflected in the data loader, as described in Algorithm 1.
@@
      (Here $\mathcal{G}$ represent the grid)
@@
$\mathcal{L}u=f\text{ in }\Omega^{d}\text{ and }\mathcal{B}u=g\text{ on }\partial\Omega^{d}$ (2)
where, $u\in\mathbb{R}^{d}$, $\mathcal{L}$ and $\mathcal{B}$ are operators acting on the interior, $\Omega$ and boundary, $\partial\Omega$ respectively. $\mathcal{B}u=g$ is the boundary constraint. The above equation can be recast as an equivalent optimization problem with the following loss function:
$\mathcal{J}(\theta)=\lambda_{r}\|\mathcal{L}\hat{u}_{\theta}-f\|_{2}^{2}+\lambda_{b}\|\mathcal{B}\hat{u}_{\theta}-g\|_{2}^{2}$ (3)
@@
$\hat{u}_{\theta}(X_{1},X_{2},\ldots,X_{B})=f_{1}^{(\theta_{1})}(X_{1})\otimes f_{2}^{(\theta_{2})}(X_{2})\otimes\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ \otimes f_{B}^{(\theta_{B})}(X_{B}).$
@@
Note that, since the PDE loss is computed only over the active dimensions, all inactive dimensions are assigned constant values based on their respective spatial locations in a computational domain.

### v3

@@ Abstract
High-dimensional partial differential equations (PDEs) arise in diverse scientific and engineering applications but remain computationally intractable due to the curse of dimensionality. Traditional numerical methods struggle with the exponential growth in computational complexity, particularly on hypercubic domains, where the number of required collocation points increases rapidly with dimensionality. Here, we introduce Anant-Net, an efficient neural surrogate that overcomes this challenge, enabling the solution of PDEs in high dimensions. Unlike hyperspheres, where the internal volume diminishes as dimensionality increases, hypercubes retain or expand their volume (for unit or larger length), making high-dimensional computations significantly more demanding. Anant-Net efficiently incorporates high-dimensional boundary conditions and minimizes the PDE residual at high-dimensional collocation points. To enhance interpretability, we integrate Kolmogorov–Arnold networks into the Anant-Net architecture. We benchmark Anant-Net’s performance on several linear and nonlinear high-dimensional equations, including the Poisson, Sine-Gordon, and Allen-Cahn equations, as well as transient heat equations, demonstrating high accuracy and robustness across randomly sampled test points from high-dimensional spaces. Importantly, Anant-Net achieves these results with remarkable efficiency, solving 300-dimensional problems on a single GPU within a few hours. We also compare Anant-Net’s results for accuracy and runtime with other state-of-the-art methods. Our findings establish Anant-Net as an accurate, interpretable, and scalable framework for efficiently solving high-dimensional PDEs. The Anant-Net code is available at https://github.com/ParamIntelligence/Anant-Net.
@@ 1 Introduction
Physics-informed deep learning (PIDL) represents a rapidly advancing framework that integrates known governing physical laws, typically formulated as PDEs, into the training process of deep neural networks. In contrast to conventional data-driven models that rely solely on observational data, PIDL incorporates physical constraints to guide learning, thereby enhancing generalization, reducing data dependence, and improving interpretability. This synthesis of physics and deep learning has demonstrated broad applicability in solving forward and inverse problems across scientific and engineering domains, particularly in scenarios involving limited, noisy, or deceptive data. Key methodologies under the PIDL umbrella include physics-informed neural networks (PINNs) [1, 2, 3, 4], which embed PDE constraints via automatic differentiation; sparse identification of nonlinear dynamics (SINDy) [5, 6], which infers governing equations by promoting sparsity in learned representations; and physics-informed neural operators [7, 8, 9, 10, 11], which approximate solution operators across function spaces to model families of PDEs. These approaches are particularly well-suited for high-dimensional problems, where traditional numerical solvers suffer from the curse of dimensionality.
High-dimensional PDEs are integral to various scientific and engineering domains, including quantum mechanics, financial mathematics, and optimal control. Their solutions provide crucial insights into complex, multi-scale phenomena that cannot be accurately captured using lower-dimensional approximations. However, solving these equations efficiently remains a significant challenge due to the curse of dimensionality, the exponential growth in computational complexity and data requirements as the number of dimensions increases. This issue is particularly pronounced in numerical approximations, making high-dimensional PDEs computationally intractable with conventional methods. In this work, we introduce Anant-Net (with Anant meaning infinite in Sanskrit). While the method does not address infinite-dimensional problems, it is designed to tackle finite but high-dimensional PDEs. We believe that Anant-Net provides an efficient and scalable approach for solving such problems. The Anant-Net is designed to mitigate the curse of dimensionality and enable the efficient solution of high-dimensional PDEs critical to real-world applications. Our approach provides an efficient framework for handling large datasets, whose size typically scales exponentially with dimensionality. While purely data-driven models might seem a viable alternative, they often struggle with excessive data requirements, making them impractical for high-dimensional problems. To address this, we leverage physics-informed neural surrogates, incorporating the underlying PDE as an additional constraint. This integration significantly reduces the dependency on large datasets, offering a more scalable and computationally efficient framework for solving high-dimensional PDEs.
@@ 1.1 Related Work
In the context of high-dimensional PDEs, two primary factors determine the efficacy of any method based on scientific machine learning: (1) the sampling strategy for data or collocation points, and (2) the computation of gradients for residual losses via automatic differentiation, which are used as soft constraints in training physics-informed models. It is evident that a significant number of existing methods represent a trade-off between these factors, which ultimately influences their scalability and suitability for solving real-world problems of practical relevance. While various methods exist for solving high-dimensional PDEs, this work focuses on neural surrogates that not only leverage available data but also incorporate physical laws, ensuring the consistency of the solutions with underlying governing principles. In recent years, the challenge of solving high-dimensional PDEs has been addressed by several researchers through the use of neural surrogate methods. Han et al. [12] introduced a deep neural network framework that reformulates high-dimensional PDEs as backward stochastic differential equations. However, their approach relied on approximating the gradient of the solution using a neural network, which introduced scalability issues, particularly in scenarios involving long-time integration. This limitation was subsequently addressed by Raissi et al. [13], who proposed an alternative method that directly models the solution of the PDE with a neural network, rather than its gradient. The efficacy of this improvement was demonstrated through its application to the Black-Scholes-Barenblatt and Hamilton-Jacobi-Bellman equations in 100 dimensions. Hu et al. [14] introduced a stochastic dimensions gradient descent (SDGD) framework based on physics-informed neural networks to solve high-dimensional PDEs on hypersphere. This method stochastically samples the dimensions of the PDE and computes the residuals only in the selected sample dimensions, rather than performing the computation across all dimensions. Their approach significantly reduced the computational burden on hardware, leading to notable speed-ups and improved accuracy compared to other contemporary methods for high-dimensional PDEs. However, their method bypassed the need for data or boundary point sampling by directly enforcing boundary conditions within the model, particularly for hypersphere, where the internal volume approaches zero as the number of dimensions increases (see Figure 1). In the similar line of work, Hu et al.[15] proposed Monte Carlo fractional PINN (MC-fPINN) and its extension, MC-tfPINN for tempered fractional PINN (an extension to fPINN [16]) to solve extremely high-dimensional problems. Further, an alternate approach was proposed by the same authors [17] extended fractional PINN to solve high-dimensional Fokker-Planck-Lévy (FPL) equations in high-dimensions. Adaptive training methods have also been proposed as supplementary strategies to enhance the accuracy of physics-informed neural network models for high-dimensional PDEs [18].
@@
Several studies have leveraged the weak formulation of PDEs to train deep learning models, particularly for high-dimensional problems. A pioneering approach in this direction is the Deep Ritz method [25], which utilizes the variational formulation of PDEs by minimizing an energy functional. This method, designed to address high-dimensional PDEs, employs a ResNet-based architecture and is trained using stochastic gradient descent. It systematically samples data from both the interior and the boundary of the domain, distinguishing itself from approaches such as that of Hu et al. [14], which circumvent boundary data sampling. Another method utilizes the variational formulation or the weak form of the PDE was proposed by Samaniego et al. [26]. However, their method was limited to solving low dimensional PDEs and couldn’t be extended for the cases discussed henceforth in this paper. Building on this framework, the Weak Adversarial Network (WAN) [27] was introduced to solve high-dimensional PDEs by reformulating the problem of finding a weak solution as an operator norm minimization task. In contrast to conventional numerical methods such as finite difference schemes, WAN demonstrated computational efficiency, stability, and the ability to mitigate the curse of dimensionality. The methodology employs generative adversarial network (GAN) framework, where the weak solution and test functions are parameterized by deep neural networks, termed the primal and adversarial networks, respectively. More recently, the WAN framework has been extended through the introduction of XNODE-WAN [28], which replaces the primal network with a Neural Ordinary Differential Equation (Neural ODE) model. This enhancement has resulted in improved computational speed and accuracy, particularly for time-dependent high-dimensional PDEs. These approaches systematically sample data across both the interior and the boundary of the computational domain to enhance the robustness of their solutions.
@@ 1.2 Our Contributions
  • We also introduce an interpretable variant of Anant-Net based on the Kolmogorov–Arnold Network (KAN), called as Anant-KAN, enabling further insights into the learned solution structures in high-dimensional settings. To the best of our knowledge, this is the first application of KAN to high-dimensional problems.
  • We benchmark Anant-Net against several state-of-the-art methods, reporting comparative results in terms of runtime as well as the mean and variance of relative testing error (%), and observe that Anant-Net consistently outperforms the existing methods.
The paper is organized as follows: Section 2 provides a detailed description of the Anant-Net architecture, including the methodologies employed for organizing training data and incorporating collocation points. Furthermore, it introduces the interpretable KAN-based variant of Anant-Net, referred to as Anant-KAN. Section 3 presents computational results for a representative set of high-dimensional boundary and initial value problems, encompassing both linear and nonlinear, as well as steady-state and transient cases. A comparative analysis against existing state-of-the-art architectures is also provided. To this end, Section 4 highlights the key findings of this work and provides a detailed discussion of the advantages and limitations of the proposed methodology.
@@ 2 Methodology: Anant-Net Architecture
Figure 2: Schematic representation of general Anant-Net architecture with batch-size $B$, highlighting its core components and computational flow for solving high-dimensional PDEs. We partition the $d$-dimensional input space into $B$ equal segments, each of size $d/B$. The first body network receives the subset $x_{1},x_{2},\dots,x_{\frac{d}{B}}$}. The second body network processes $x_{\frac{d}{B}+1},\dots,x_{\frac{2d}{B}}$}, while the third network processess $x_{\frac{2d}{B}+1},\dots,x_{\frac{3d}{B}}$}, and so on. The final (i.e., $B^{\text{th}}$) body network is assigned the remaining inputs $x_{\frac{(B-1)d}{B}+1},\dots,x_{d}$}.
@@
Figure 2 shows the general architecture of Anant-Net. The architecture comprises a number of body networks equal to the batch size. Let $u$ denotes the solution of a $d$-dimensional PDE, which is the unknown to be approximated by Anant-Net in a $d$-dimensional hypercube. For brevity, we consider a batch size of three, the Anant-Net prediction of the solution field, denoted by $\hat{u}_{\theta}$ , is expressed as:
$\hat{u}_{\theta}(X_{1},X_{2},X_{3})=f^{(\theta_{1})}_{1}(X_{1})\otimes f^{(\theta_{2})}_{2}(X_{2})\otimes f^{(\theta_{3})}_{3}(X_{3})=\sum_{j}^{emb}f^{(\theta_{1})}_{1,j}(X_{1})f^{(\theta_{2})}_{2,j}(X_{2})f^{(\theta_{3})}_{3,j}(X_{3}),$ (1)
where $\otimes$ represents the tensor product notation and the inputs are partitioned as $X_{1}=\{x_{1},\dots,x_{d/3}\}$, $X_{2}=\{x_{d/3+1},\dots,x_{2d/3}\}$ , and $X_{3}=\{x_{2d/3+1},\dots,x_{d}\}$, such that $X_{1}\cap X_{2}\cap X_{3}=\emptyset$. The indices $p\in\mathrm{D}_{1}=\{1,2,\ldots,d/3\},q\in\mathrm{D}_{2}=\{d/3+1,2,\ldots,2d/3\},r\in\mathrm{D}_{3}=\{2d/3+1,2,\ldots,d\}$ , corresponding to the number of batch elements, are stochastically sampled from the set of active dimensions, i.e., $\{x_{p},x_{q},x_{r}\}\subset X=X_{1}\cup X_{2}\cup X_{3}\in\mathbb{R}^{d}$, which sweeps over all dimensional indices. The network parameters are given by $\theta=\bigcup\theta_{i},\ \forall i$ (all body networks), where $f_{i}^{(\theta_{i})}$ denotes the output vector of the $i^{\text{th}}$ body network (a feed-forward neural network), which can be given as
$f_{i}^{(\theta_{i})}(X_{i})=\Psi_{L}\circ\Psi_{L-1}\circ\dots\circ\Psi_{1}(X_{i}),$
where $\Psi_{\ell}(Z)=\sigma_{\ell}(W_{\ell}Z+b_{\ell})$, with $\sigma_{\ell}$, $W_{\ell}$, and $b_{\ell}$ denoting the activation function, weight matrix, and bias vector at layer $\ell$, respectively, and $L$ is the total number of hidden-layers. Here $f^{(\theta_{i})}_{i,j}$ represents $j^{\text{th}}$ component of the embedding vector $f^{(\theta_{i})}_{i}\in\mathbb{R}^{emb}$. Although there is no strict ordering required for partitioning the dimension set $\mathrm{D}=\mathrm{D}_{1}\cup\mathrm{D}_{2}\cup\mathrm{D}_{3}$, we adopt an ascending-order partitioning strategy for all experiments presented in this paper for the sake of consistency. The number of partitioned subsets is arbitrary and can be increased by proportionally increasing the number of body networks in the Anant-Net architecture. However, for brevity, we fix the total number of partitioned sets; and correspondingly, the number of body networks, to three in all experiments discussed herein. Algorithm 1 gives the details of training data and collocation points for the Anant-Net architecture. It is also important to note that modifying the number of body networks by changing the number of partitioned dimension subsets must be reflected in the data loader, as described in Algorithm 1.
@@
  (Here $\mathcal{G}_{data}^{pqr}$ represents grid of data points and $\mathcal{G}_{coll}^{pqr}$ represents grid of collocation points)
@@
$\mathcal{L}u=f\text{ in }\Omega^{d}\text{ and }\mathcal{B}u=g\text{ on }\partial\Omega^{d},$ (2)
where, $u$ denotes the solution of a $d$-dimensional PDE, which is the unknown, $\mathcal{L}$ and $\mathcal{B}$ are operators acting on the interior, $\Omega$ and boundary, $\partial\Omega$ respectively. $\mathcal{B}u=g$ is the boundary constraint. The above equation can be recast as an equivalent optimization problem with the following loss function:
$\mathcal{J}(\theta)=\lambda_{r}\|\mathcal{L}\hat{u}_{\theta}-f\|_{2}^{2}+\lambda_{b}\|\mathcal{B}\hat{u}_{\theta}-g\|_{2}^{2},$ (3)
@@
$\hat{u}_{\theta}(X_{1},X_{2},\ldots,X_{B})=f^{(\theta_{1})}_{1}(X_{1})\otimes f^{(\theta_{2})}_{2}(X_{2})\otimes\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ \otimes f^{(\theta_{B})}_{B}(X_{B}).$
@@
Note that, since the PDE loss is computed only over the active dimensions, all inactive dimensions do not contribute towards the gradient computation since they represent fixed locations in the computational domain.